OpenStudy (anonymous):

{n^2/2^n} does the sequence converge or diverge if it converges how do u find the limit?

OpenStudy (anonymous):

Use the ratio test: \[\lim_{n \rightarrow \infty} \frac{\frac{(n+1)^2}{2^{n+1}}}{\frac{2^n}{n^2}}\]

OpenStudy (anonymous):

= \[\lim_{n \rightarrow \infty} \frac{(n+1)^2}{2^{n+1}} \frac{2^n}{n^2} = \lim_{n \rightarrow \infty} \frac{1}{2}\frac{(n+1)^2}{n^2}\]

OpenStudy (anonymous):

or you could watch the vid: http://www.youtube.com/watch?v=wV1FrqwZyKw&feature=feedlik

OpenStudy (anonymous):

which equals (1/2). So the ratio test says that if the ratio yeilds something less than 1, the series converges

OpenStudy (anonymous):

ohh ok i was trying to aproach this using the nth term test

OpenStudy (anonymous):

if you try using the n-th term test, you'll get 0, which is inconclusive.